Circular No. 9006

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V1647 ORIONIS D. K. Ojha, S. K. Ghosh, and S. S. Kaurav, Tata Institute of Funda-

mental Research, Mumbai; B. C. Bhatt and D. K. Sahu, Indian Institute of Astrophysics, Bangalore; and A. Tej, Indian Institute of Space Science and Technology, Thiruvananthapuram, report the following optical and nearinfrared magnitudes of the young eruptive star V1647 Ori (cf. IAUCs 8284, 8354, 8396, 8600, 8681, 8694), made with the 2-m Himalayan Chandra telescope (+ HFOSC spectrograph and NIR camera) at Hanle (Ladakh) and the 2-m IUCAA Girawali telescope (+ IFOSC spectrograph) at Girawali (Pune): Sept. 14 UT, V = 18.92 ± 0.03, R = 17.03 ± 0.02, I = 14.91 ± 0.01; Oct. 19, J = 10.64 ± 0.01, H = 9.01 ± 0.01, K = 7.63 ± 0.01 (measured in a 7′′.2-radius aperture); Nov. 3, V = 18.91 ± 0.03, R = 17.09 ± 0.02. No significant variation in brightness of V1647 Ori was seen during the period of about two months since the recent new outburst phase (cf. IAUCs 8968, 8969). In comparison with the last reported quiescent phase (cf. Ojha et al. 2006, MNRAS 368, 825), there has been a brightening of ∼ 3 magnitudes in R, and the infrared colors suggest that circumstellar matter of A_{v} ∼ 7.5 mag has probably been cleared during this recent outburst. Flux-calibrated optical spectra of V1647 Ori obtained from Sept. 14–Dec. 5 show strong Hα emission with blue-shifted P-Cyg absorption and the Ca II triplet lines in emission. The dip in the P-Cyg profile appears to occur around 655.2 nm, which indicates an in-fall velocity of the order of ∼ 500 km/s. The equivalent widths of the Hα and Ca II lines are ∼ −2.5 and −1.0 nm, respectively, which are similar to the values seen in the previous outburst (Ojha et al., op.cit.). Regular monitoring of V1647 Ori is underway in optical and near-infrared wavebands.

COMET C/2008 R7 (SOHO) Further to IAUC 9003, an additional near-sun Kracht-group presumed

comet was found by R. Kracht on SOHO website images (“discovery” image tabulated below). C/2008 R7 was stellar in appearance in C3, C2, and COR2-A images, peaking at mag ∼ 5.5 around Sept. 6.83 UT. Kracht suggests that C/2008 R7 may be identical with C/2002 S5, though B. G. Marsden notes that identity with C/2002 S11 is also possible (despite its proposed association with C/2008 G6).

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 R7 Sept. 6.447 10^{h}52.^{m}3 + 7^{o}37′ C3/2* RK 2008-S67
2008 December 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green